The perimeter of a square is 12 meters greater than the perimeter of another square. It's area is 45 square meters greater than the area of the smaller square. Find the perimeter of each square. : The perimeter of a square is 12 meters greater than the perimeter of another square. It's area is 45 square meters greater than the area of the smaller square. Find the perimeter of each square.
Answer by nerdybill(1120)

The perimeter of a square is 12 meters greater than the perimeter of another square. It's area is 45 square meters greater than the area of the smaller square. Find the perimeter of each square.
.
Let x = perimeter of smaller square
then from: "12 meters greater than the perimeter of another square"
x+12 = perimeter of larger square
.
length of one side of smaller square = x/4
length of one side of larger square = (x+12)/4
.
(x/4)^2 + 45 = ((x+12)/4)^2
(x^2/16) + 45 = (x^2+24x+144)/16
x^2 + 720 = x^2+24x+144
720 = 24x+144
720 = 24x+144
576 = 24x
24 meters = x (perimeter of smaller square)
.
x+12 = 24+12 = 36 meters (perimeter of larger square)